# Rapid near-inertial wave generation at topography in response to surface wind forcing

### v2

@@ 5 Discussion and conclusions
As hypothesised above, the topographic near-inertial waves may contribute to deep ocean mixing, but the exact extent and location of enhanced bottom mixing due to our proposed mechanism remains an open question. Our study was designed to maximise both wave generation and propagation, and hence mechanisms for wave breaking and dissipation were intentionally either removed or suppressed. It remains possible that in more realistic conditions the waves would immediately break, become trapped, or otherwise dissipate close to the generation site. Further study is therefore needed to understand how these waves would behave in the real ocean. If wind-forced topographic near-inertial waves are responsible for intensified near-bottom mixing, then one could assess their impact indirectly by measuring the bottom mixing rates at other intersection points of topographic features and storm systems. This could be done either with existing moorings, or by analysing model outputs of high resolution ocean models with more realistic forcing and bathymetry. The advantage of using a high-resolution ocean models is that the effects of tides could easily be removed, so enhanced internal wave generation or breaking at the bottom in the presence of storms could more easily be teased out. It would also be worthwhile to apply these analysis techniques to the data of the previously mentioned studies (van Haren, 2020; Morozov and Velarde, 2008; Jouanno et al., 2016) which found evidence for bottom NIW activity immediately after a storm, to investigate whether this mechanism is indeed consistent with these observations.
@@
In conclusion, we have studied a novel internal-wave generation mechanism whereby the rapid de-phasing of a wind-driven internal waves modes results in bottom flow over topographic features. The energy flux associated with these wind-driven topographic near-inertial waves reached $0.2\;\mathrm{mW}\,\mathrm{m}^{-2}$, a value $10\%$ the size of surface generated NIWs in the same domain. By separating the topographic near-inertial waves from other dynamical features in our idealised experiments as cleanly as possible, we found that wave energy scales in a way consistent with existing internal wave theory. Further study of topographic near-inertial waves in ocean observations and realistic models would help us clarify how these waves might fit into the broader internal wave spectrum, and our picture of abyssal mixing.

### v4

@@ 5 Discussion and conclusions
The results of our idealised study suggest that the scatter of storm-generated NIWs at topography may contribute to energy downscaling and in turn deep ocean mixing, but the exact extent and location of enhanced bottom mixing due to our proposed mechanism remains an open question. It remains possible that in more realistic conditions the waves would immediately break, become trapped, or otherwise dissipate close to the generation site, which was intentionally avoided in our study to better measure the energy fluxes. Further study is therefore needed to understand how these waves would behave in the real ocean. If wind-forced topographic near-inertial waves are responsible for energy downscaling and intensified near-bottom mixing, then one could assess their impact indirectly by measuring the wavenumber spectra and bottom mixing rates at other intersection points of topographic features and storm systems. This could be done either with existing moorings, or by analysing model outputs of high resolution ocean models with more realistic forcing and bathymetry. The advantage of using high-resolution ocean models is that the effects of tides could easily be removed, so enhanced internal wave generation or breaking at the bottom in the presence of storms could more easily be teased out. It would also be worthwhile to apply these analysis techniques to the data of the previously mentioned studies (van Haren, 2020; Morozov and Velarde, 2008; Jouanno et al., 2016) which found evidence for bottom NIW activity immediately after a storm, to investigate whether this mechanism is indeed consistent with these observations.
@@
In conclusion, we have studied a novel internal-wave generation mechanism whereby the rapid dephasing of a wind-driven internal waves modes results in bottom flow over topographic features. The energy flux associated with these wind-driven topographic near-inertial waves reached $0.2\;\mathrm{mW}\,\mathrm{m}^{-2}$, a value $10\%$ the size of surface-generated NIWs in the same domain. By separating the topographic near-inertial waves from other dynamical features in our idealised experiments, we found that wave energy scales in a way consistent with existing internal wave theory. Further study of topographic near-inertial waves in ocean observations and realistic models would help us clarify how these waves might fit into the broader internal wave spectrum, and our picture of abyssal mixing.